BROADBAND CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE SPECTROSCOPY AND MOLECULAR STRUCTURE OF THE ARGON-1-CHLORO-1-FLUOROETHYLENE COMPLEX

\underline{MARK D. MARSHALL} AND HELEN O. LEUNG, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

Previous studies of argon complexes with fluoroethylenes have revealed a preference for a geometry that maximizes the contact of the argon atom with heavy atoms on the fluoroethylene.^{a} We have observed a continuation of this trend when one of the fluorine atoms is replaced by chlorine. As part of a systematic study of the effect of chlorine substitution on intermolecular interactions, we have examined the argon-1-chloro-1-fluoroethylene complex, and obtained the 5.6 – 18.1 GHz chirped-pulse Fourier transform microwave spectrum of this species. Transitions for both the ^{35}Cl and ^{37}Cl isotopologues are observed and analyzed to provide geometric parameters for this non-planar complex. The structure is found to be similar to those of analogous complexes and agrees well with ab initio predictions.

aZ. Kisiel, P.W. Fowler, and A.C. Legon, J. Chem. Phys. 95, 2283 (1991).